OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −6(2u − 1)

OpenStudy (anonymous):

-6(2u-1) -6*2=-12u -6*-1= +6 -12u+6

OpenStudy (anonymous):

thanks : )

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Identify which distributive property problem below would help you solve the following scenario: Your English class has 34 students. You would like to treat the class to cookies. You would like each person to have 3 cookies. Which expression would help you figure out how many cookies you would need to bake?

OpenStudy (anonymous):

x would be number of cookies you need. so the equation would have to equal x

OpenStudy (anonymous):

since there is 34 students and three cookies for each student. x= 3 (34)